OpenStudy (anonymous):

Use the substitution method to solve the following system of equations.4x - y = 37x - 9y = -2 A: (1, 1) B: (6, -3) C: (6, 1) D: (1, -3) *I can help you with Psych, English, and Philosophy in return!

OpenStudy (anonymous):

So, is the system\[4x-y=37\]\[x-9y=-2\]?

OpenStudy (anonymous):

I don't know what the system is

OpenStudy (anonymous):

It asks to use the substitution method

OpenStudy (anonymous):

Steps for Using the Substitution Method in order to Solve Systems of Equations Solve 1 equation for 1 variable. (Put in y = or x = form) Substitute this expression into the other equation and solve for the missing variable. Substitute your answer into the first equation and solve. Check the solution.

OpenStudy (anonymous):

If so, just rewrite the bottom equation in terms of x, \[x=9y-2\] and substitute 9y-2 for x in the top equation to find y

OpenStudy (anonymous):

4(9y-2)-9=37

OpenStudy (anonymous):

I understand that much, but when it comes to finding two digits as a result, that's where I'm lost

OpenStudy (anonymous):

All right. I can show you that. What you are looking for is an ordered pair, or something like this (x,y). You can first solve for y using the equation I just gave you:\[36y-8-9=37\]\[36y=54\]\[y=\frac{ 54 }{ 36}=\frac{ 3 }{ 2 }\]Now that you have a y value, substitute it one of your original equations, say x-9y=-2\[x=9y-2=9(\frac{ 3 }{ 2 })-2=\frac{ 27 }{ 2 }-\frac{ 4 }{ 2 }\]\[x=\frac{ 23 }{ 2 }\]That's how it's done, but I think there's something wrong...? Are the answers right, or is the system of equations not what I supposed?